Assume that the sum of 100 integers n1, n2, .... is greater than 100. So

A. All the numbers n1, n2, .... have to be greater than or equal to 1.

B. If one of the numbers n1, n2, .... is lower or equal to -100, so at least one of the remaining numbers must be greater than 2

C. If between the numbers Solutions of the inequality [(4x2-3x-1)^1/2]>= 2x-3 are?



1
Expert's answer
2021-08-04T14:34:57-0400

A. False. Counterexample: let "n_1=-2<1, n_2=5, n_3=n_4=...=n_{100}=1"


"S_{100}=-2+5+1+1+...+1=3+98(1)=101>100"



B. True. Let "n_1=-100, n_2\leq 2, n_3\leq2,..., n_{100}\leq 2"


"S_{100}\leq-100+99(2)=98<100"

We have contradiction.

Therefore If one of the numbers n1, n2, .... is lower or equal to -100, so at least one of the remaining numbers must be greater than 2.


C.


"(4x^2-3x-1)^{1/2}\geq2x-3"

"4x^2-3x-1\geq0=>(x-2)(4x+1)\geq0"

"x\leq-0.25\ or\ x\geq2"

If "x\leq-0.25," then "2x-3<0."


If "x\geq2," then "2x-3>0"


"4x^2-3x-1\geq(2x-3)^2"

"4x^2-3x-1\geq4x^2-12x+9"

"9x\geq10"

"x\geq \dfrac{10}{9}"

Solution of the inequality


"x\in(-\infin, -0.25]\cup [2,\infin )"